A Bibliography of Supercomputing ’2002

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 October 2010 Version 1.03

Abstract

This bibliography records articles presented at the Supercomputing ’2002 conference.

Title word cross-reference

14.9 [SMSY02]. 16.4 [YIU^{+}02].

240-Processor [WWF02]. 26.58 [STF^{+}02]. 29.5 [MKFD02].

6 [MKFD02].

Accelerating [SLMW02]. Accelerator [SWK^{+}02]. ACM [IEE02]. Active [TCH02]. Advanced [SWK^{+}02]. Aggregate [HNA02]. Algorithm [OOS02]. AlphaServer [Wor02]. Analysis [DEHS02].

Application [dLR02]. Applications [VY02]. Applying [AZV^{+}02]. Approach [BBC^{+}02a]. Architecture [PTV02, LMN02]. Architectures [HD02]. Archives [CG02]. Array [PH02]. Arrays [CG02]. Artificial [KHLL02]. Asserting [VW02]. Atmospheric [STF^{+}02]. Atomistic [KKN^{+}02]. Automated [TCH02]. Awareness [PTV02].

Backbone [HIT^{+}02]. Baltimore [IEE02]. bandwidth [EG02]. Based [PcC02]. Beowulf [WWF02]. Better [PH02]. between [PTV02]. Biomolecular [PZKK02]. BlueGene [AAA^{+}02]. BlueGene/L [AAA^{+}02]. Bounds [VDY^{+}02]. Bridge [PTV02].

Cache [OOS02]. Cache-to-Cache [AGGD02]. Calculations [SLMW02]. Capable [PcC02]. Case [Wor02]. cc [AGGD02]. cc-NUMA [AGGD02]. CFD [DK02]. Chemistry [BBC^{+}02a]. Chimera [AZV^{+}02]. Clint [EG02]. Cluster [HD02]. Cluster-Based [PcC02]. Clusters [ASTK02]. Codes [dLR02]. Collaborative [KKN^{+}02]. Common [MAG^{+}02]. Communication [RH02]. Compact [dLR02]. Compiling [PH02]. Computational [PFY^{+}02]. Computing [BBH^{+}02]. Concepts [AZV^{+}02]. Conducting [HAS02]. Conference [IEE02]. Constructing [CDF^{+}02]. Contraction [PH02]. Counter [AV02]. Cubic [WWF02].

Daemon [DMT02]. Data [OOS02]. Data-Intensive [HIT^{+}02]. Data-Mapping [TRN02]. Decoupled [DCB02]. Density [WWF02]. Departmental [KHLL02]. Deterministic [DK02]. Development [DCB02]. dimensional [SMSY02]. Direct [YIU^{+}02]. Directory [BWS02]. Discovery [Hos02]. Disk [OOS02]. Disks [CG02]. Distributed [dLR02]. Dual [DK02]. Dual-Level [DK02]. Dynamic [MS02].

Early [WDF^{+}02]. Earth [YIU^{+}02]. Effects [HNA02]. Efficient [RH02]. Electrically [HAS02]. Empirical [VY02]. Environment [DCB02]. Environments [LGST02]. Equality [SLMW02]. Evaluation [PcC02]. Executing [SFB^{+}02]. Execution [AKSS02]. Expectations [VW02]. Experimental [BBH^{+}02].

Factorization [GL02]. Fast [HAS02]. Fault [BBC^{+}02b]. Finding [AZV^{+}02]. Fire [SJ02]. Flows [HNA02]. Fluid [SMSY02]. Forecasting [SW02]. Fourier [YIU^{+}02]. Framework [SCW^{+}02]. Fusion [SMSY02].

G [AKSS02]. GAMESS [BGE^{+}02]. Gap [PTV02]. Gauss [ABG02]. Gigabit [HIT^{+}02]. Giggle [CDF^{+}02]. Gilgamesh [SZ02]. Global [STF^{+}02]. GrADS [DCB02]. GRAPE [MKFD02]. GRAPE-6 [MKFD02]. Grid [LMN02, KHLL02]. Grids [OOS02]. Guide [DEHS02].

Hand [WCHS02]. Harmony [TCH02]. Hash [MS02]. Hierarchical [RR02]. High [BBH^{+}02]. High-bandwidth [EG02]. High-Density [WWF02]. High-Level [BBC^{+}02a]. High-Performance [BBC^{+}02a]. HPF [SMSY02].

IBM [MS02]. ICENI [LMN02].

Idle [CG02]. Implementation [PcC02]. Implemented [LMN02]. Implementing [Tra02]. Improving [GcC02]. Increasing [PTV02]. Infrastructure [DEHS02]. Insight [IEE02]. Inspection [KHLL02]. Instrumentation [NZ02]. Integration [BGE^{+}02]. Intelligent [KHLL02]. Intensive [HIT^{+}02]. Interconnect [SJ02]. Interfaces [ASTK02]. Interoperable [PFY^{+}02]. Inverse [ABG02]. IP [PcC02].

Japan [KKN^{+}02]. Jini [LMN02].

Keys [MAG^{+}02]. Krylov [ABG02].

L [AAA^{+}02]. LAPI [MS02].

latency [EG02].

Latency-Tolerant [TRN02]. Level [BBC^{+}02a]. Library [RR02, CGG02]. Lightning [FPF^{+}02]. Lightning-Fast [FPF^{+}02]. Likelihood [SLMW02]. Likelihood-based [SLMW02]. Link [SJ02]. Location [CDF^{+}02]. Lookup [GcC02]. Loops [ASTK02]. Loss [HNA02]. Low [EG02]. Low-latency [EG02]. LU [GL02].

Management [FPF^{+}02]. Managers [OOS02]. Mapped [ASTK02]. Mapping [TRN02]. Massive [CG02]. Mathematics [BBH^{+}02]. Matrix [VDY^{+}02]. Matrix-Multiply [PTV02]. Matrix-Vector [VDY^{+}02]. Maximum [SLMW02]. MD [IEE02]. Mechanical/Classical [KKN^{+}02]. Mechanics [BPR^{+}02]. Mechanism [Tra02]. Meets [BBH^{+}02]. Memory [DEHS02]. Merging [MAG^{+}02]. Message [CGG02]. Meter [WWF02]. Method [HAS02]. Methods [ABG02]. Metrics [AV02]. Microprocessor [AV02]. Misses [AGGD02]. Modeling [SCW^{+}02]. Monitoring [LGST02]. MPI [HD02]. MPICH [BBC^{+}02b]. MPICH-V [BBC^{+}02b]. Multi [HIT^{+}02]. Multi-Gigabit [HIT^{+}02]. Multi-Processor [RR02]. Multiple [SFB^{+}02]. Multiply [VDY^{+}02]. Multipole [HAS02]. Multiprotocol [CGG02]. Multiscale [ABG02]. Multithreaded [SZ02]. Multivariate [SW02].

NAMD [PZKK02]. Neptune [MKFD02]. Nested [ASTK02]. Network [GcC02, SW02]. Networks [MAG^{+}02]. Newton [ABG02]. Nodes [BBC^{+}02b]. Nonuniform [RH02].

November [IEE02]. NPB [EGC02].

NUMA [AGGD02]. Numerical [YIU^{+}02].

One [WWF02]. onto [ASTK02]. Open [LMN02]. OpenMP [HD02]. Operations [SFB^{+}02]. Optimizations [VDY^{+}02]. Optimizing [AKSS02]. Overview [AAA^{+}02]. Owner [AGGD02].

p690 [WDF^{+}02]. Packet [HNA02]. Paradigms [HD02]. Parallel [dLR02]. Parallelism [DK02]. Particle [SWK^{+}02]. PC [KKN^{+}02]. Peak [PTV02]. Classical [KKN^{+}02]. Perfect [HAS02]. Performance [AV02]. Personalized [WCHS02]. Petaflops [SZ02]. Phylogenetic [SLMW02]. Pipelined [ASTK02]. Pivoting [GL02]. planetesimals [MKFD02]. Portals [WCHS02]. Potential [EGC02]. Prediction [SCW^{+}02]. Proactivity [BWS02]. Problems [DK02]. Proceedings [IEE02]. Process [Tra02]. Processor [GcC02]. Processor-In-Memory [SZ02]. Processors [PZKK02]. Program [PTV02]. Programs [PH02]. Propagation [ABG02]. Proteus [CGG02]. Proxy [AKSS02]. Proxy-G [AKSS02].

QMView [BGE^{+}02]. QoS [PcC02].

QoS-Capable [PcC02]. Quantum [BBC^{+}02a]. Quantum-Mechanical [KKN^{+}02]. Quantum-Mechanical/ Classical [KKN^{+}02]. Query [AKSS02].

region [MKFD02]. Replacement [OOS02]. Replica [CDF^{+}02]. Research [HIT^{+}02]. Reservoir [HIT^{+}02]. Resource [SW02]. Revisited [PTV02]. Route [GcC02]. Router [PcC02].

Salinas [BPR^{+}02]. SC2002 [IEE02]. Scalable [AV02]. Scaling [Wor02]. Scattering [HAS02]. Scheduling [ASTK02]. Scheme [TRN02]. Science [SMSY02]. Scientific [dLR02]. Separated [EG02]. Service [LMN02, SW02]. Services [PFY^{+}02]. SIGMA [DEHS02]. Signatures [dLR02]. simulation [MKFD02]. Simulations [SWK^{+}02]. Simulator [DEHS02, YIU^{+}02]. Sky [AZV^{+}02]. Sloan [AZV^{+}02]. SmartPointers [WCHS02]. SMP [HD02]. SMPs [ASTK02]. Software [BPR^{+}02]. Solid [BPR^{+}02]. Solution [TRN02]. Sound [KHLL02]. Sparse [GL02]. Spectral [YIU^{+}02]. Static [GL02]. Stochastic [DK02]. Storage [OOS02]. STORM [FPF^{+}02]. Streams [MAG^{+}02]. Structural [BPR^{+}02]. Study [EGC02]. Subtree [SLMW02]. Sun [SJ02]. Supercomputer [AAA^{+}02]. Support [RR02]. Surfaces [HAS02]. Survey [AZV^{+}02]. Sustained [PTV02]. Synchronization [RH02]. Synthesis [BBC^{+}02a]. System [NZ02]. Systemic [HNA02].

Tables [MS02]. Tapping [KHLL02]. Tasks [RR02]. TCP [HNA02]. Techniques [AV02]. Technology [SWK^{+}02]. Terabytes [IEE02]. Terascale [SWK^{+}02]. Tflops [MKFD02]. Thousands [PZKK02]. Three [SMSY02]. Three-dimensional [SMSY02]. Tiled [ASTK02]. Tiling [PH02]. Tolerant [TRN02]. Topology [Tra02]. Transfer [AGGD02]. Transform [STF^{+}02]. Tree [SLMW02]. Triangular [TRN02]. Tuning [DMT02]. Turbulence [YIU^{+}02].

Ultra [SJ02]. Ultra-High [SJ02]. Unscalable [Wor02]. UPC [EGC02].

Uranus [MKFD02].

Uranus-Neptune [MKFD02]. USA [IEE02]. Using [MS02]. Utilization [HIT^{+}02].

V [BBC^{+}02b]. Vector [VDY^{+}02].

Vectors [SLMW02]. Virtual [AZV^{+}02]. Visualization [SWK^{+}02]. Volatile [BBC^{+}02b].

Wave [ABG02]. Weather [SW02]. Web [Hos02]. Wide [BGE^{+}02]. World [BGE^{+}02].

References

Adiga:2002:OBS

[AAA^{+}02] N. R. Adiga, G. Almasi, G. S. Almasi, Y. Aridor, R. Barik, D. Beece, R. Bellofatto, G. Bhanot, R. Bickford, M. Blumrich, A. A. Bright, J. Brunheroto, C. Cacaval, J. Castaños, W. Chan, L. Ceze, P. Coteus, S. Chatterjee, D. Chen, G. Chiu, T. M. Cipolla, P. Crumley, K. M. Desai, A. Deutsch, T. Domany, M. B. Dombrowa, W. Donath, M. Eleftheriou, C. Erway, J. Esch, B. Fitch, J. Gagliano, A. Gara, R. Garg, R. Germain, M. E. Giampapa, B. Gopalsamy, J. Gunnels, M. Gupta, F. Gustavson, S. Hall, R. A. Haring, D. Heidel, P. Heidelberger, L. M. Herger, D. Hoenicke, R. D. Jackson, T. Jamal-Eddine, G. V. Kopcsay, E. Krevat, M. P. Kurhekar, A. P. Lanzetta, D. Lieber, L. K. Liu, M. Lu, M. Mendell, A. Misra, Y. Moatti, L. Mok, J. E. Moreira, B. J. Nathanson, M. Newton, M. Ohmacht, A. Oliner, V. Pandit, R. B. Pudota, R. Rand, R. Regan, B. Rubin, A. Ruehli, S. Rus, R. K. Sahoo, A. Sanomiya, E. Schenfeld, M. Sharma, E. Shmueli, S. Singh, P. Song, V. Srinivasan, B. D. Steinmacher-Burow, K. Strauss, C. Surovic, R. Swetz, T. Takken, R. B. Tremaine, M. Tsao, A. R. Umamaheshwaran, P. Verma, P. Vranas, T. J. C. Ward, M. Wazlowski, W. Barrett, C. Engel, B. Drehmel, B. Hilgart, D. Hill, F. Kasemkhani, D. Krolak, C. T. Li, T. Liebsch, J. Marcella, A. Muff, A. Okomo, M. Rouse, A. Schram, M. Tubbs, G. Ulsh, C. Wait, J. Wittrup, M. Bae, K. Dockser, L. Kissel, M. K. Seager, J. S. Vetter, and K. Yates. An overview of the BlueGene/L supercomputer. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap207.pdf.

Akcelik:2002:PMG

[ABG02] Volkan Akcelik, George Biros, and Omar Ghattas. Parallel multiscale Gauss–Newton–Krylov methods for inverse wave propagation. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap330.pdf.

Acacio:2002:OPA

[AGGD02] Manuel E. Acacio, Jose Gonzalez, Jose M. Garcia, and Jose Duato. Owner prediction for accelerating cache-to-cache transfer misses in a cc-NUMA architecture. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap102.pdf.

Andrade:2002:APG

[AKSS02] Henrique Andrade, Tahsin Kurc, Alan Sussman, and Joel Saltz. Active Proxy-G: Optimizing the query execution process in the grid. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap219.pdf.

Athanasaki:2002:PST

[ASTK02] Maria Athanasaki, Aristidis Sotiropoulos, Georgios Tsoukalas, and Nectarios Koziris. Pipelined scheduling of tiled nested loops onto clusters of SMPs using memory mapped network interfaces. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap132.pdf.

Ahn:2002:SAT

[AV02] Dong H. Ahn and Jeffrey S. Vetter. Scalable analysis techniques for microprocessor performance counter metrics. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap257.pdf.

Annis:2002:ACV

[AZV^{+}02] James Annis, Yong Zhao, Jens Voeckler, Michael Wilde, Steve Kent, and Ian Foster. Applying Chimera virtual data concepts to cluster finding in the Sloan Sky Survey. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap299.pdf.

Baumgartner:2002:HLA

[BBC^{+}02a] Gerald Baumgartner, David E.

Bernholdt, Daniel Cociorva, Robert Harrison, So Hirata, Chi-Chung Lam, Marcel Nooijen, Russell Pitzer, J. Ramanujam, and P. Sadayappan. A highlevel approach to synthesis of high-performance codes for quantum chemistry. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap306.pdf.

Bosilca:2002:MVT

[BBC^{+}02b] George Bosilca, Aurelien Bouteiller, Franck Cappello, Samir Djilali, Gilles Fedak, Cecile Germain, Thomas Herault, Pierre Lemarinier, Oleg Lodygensky, Frederic Magniette, Vincent Neri, and Anton Selikhov. MPICH-V: Toward a scalable fault tolerant MPI for volatile nodes. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap298.pdf.

Bailey:2002:HPC

[BBH^{+}02] David H. Bailey, David Broadhurst, Yozo Hida, Xiaoye S. Li, and Brandon Thompson. High performance computing meets experimental mathematics. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap124.pdf.

Baldridge:2002:QGI

[BGE^{+}02] Kim K. Baldridge, Jerry P. Greenberg, Stephen T. Elbert, Stephen Mock, and Philip Papadopoulos. QMView and GAMESS: Integration into the World Wide Computational Grid. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap141.pdf.

Bhardwaj:2002:SSS

[BPR^{+}02] Manoj Bhardwaj, Kendall Pierson, Garth Reese, Tim Walsh, David Day, Ken Alvin, James Peery, Charbel Farhat, and Michel Lesoinne. Salinas: A scalable software for highperformance structural and solid mechanics simulations. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

                                        /www.sc-2002.org/paperpdfs/
                                        pap.pap216.pdf.

Bustamante:2002:SDS

[BWS02] Fabian E. Bustamante, Patrick Widener, and Karsten Schwan. Scalable directory services using proactivity. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

                                        /www.sc-2002.org/paperpdfs/
                                        pap.pap202.pdf.

Chervenak:2002:GFC

[CDF^{+}02] Ann Chervenak, Ewa Deelman, Ian Foster, Leanne Guy, Wolfgang Hoschek, Adriana Iamnitchi, Carl Kesselman, Peter Kunszt, Matei Ripeanu, Bob Schwartzkopf, Heinz Stockinger, Kurt Stockinger, and Brian Tierney. Giggle: A framework for constructing scalable replica location services. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap239.pdf.

Colarelli:2002:MAI

[CG02] Dennis Colarelli and Dirk Grunwald. Massive arrays of idle disks for storage archives. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap312.pdf.

Chiu:2002:PMM

[CGG02] Kenneth Chiu, Madhusudhan Govindaraju, and Dennis Gannon. The Proteus Multiprotocol Message Library. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap315.pdf.

Dail:2002:DSA

[DCB02] Holly Dail, Henri Casanova, and Fran Berman. A decoupled scheduling approach for the GrADS program development environment. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap229.pdf.

DeRose:2002:SSI

[DEHS02] Luiz DeRose, K. Ekanadham, Jeffrey Hollingsworth, and Simone Sbaraglia. SIGMA: A simulator infrastructure to guide memory analysis. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap191.pdf.

Karniadakis:2002:DLP

[DK02] Suchuan Dong and George Em. Karniadakis. Dual-level parallelism for deterministic and stochastic CFD problems. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap137.pdf.

Lu:2002:CAS

[dLR02] Charng da Lu and Daniel A. Reed. Compact application signatures for parallel and distributed scientific codes. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap244.pdf.

Dunigan:2002:TTD

[DMT02] Tom Dunigan, Matt Mathis, and Brian Tierney. A TCP tuning daemon. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

                                        /www.sc-2002.org/paperpdfs/
                                        pap.pap151.pdf.

Eberle:2002:SHB

[EG02] Hans Eberle and Nils Gura. Separated high-bandwidth and lowlatency communication in the cluster interconnect Clint. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap259.pdf.

El-Ghazawi:2002:UPP

[EGC02] Tarek El-Ghazawi and François Cantonnet. UPC performance and potential: A NPB experimental study. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap316.pdf.

Frachtenberg:2002:SLF

[FPF^{+}02] Eitan Frachtenberg, Fabrizio Petrini, Juan Fernandez, Salvador Coll, and Scott Pakin. STORM: Lightning-fast resource management. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap297.pdf.

Gopalan:2002:IRL

[GcC02] Kartik Gopalan and Tzi cker Chiueh. Improving route lookup performance using network processor cache. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap272.pdf.

Li:2002:NSA

[GL02] Laura Grigori and Xiaoye S.

Li. A new scheduling algorithm for parallel sparse LU factorization with static pivoting. In IEEE [IEE02], page ??

ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap139.pdf.

Hariharan:2002:SPF

[HAS02] Bhanu Hariharan, Srinivas Aluru, and Balasubramaniam Shanker. A scalable parallel fast multipole method for analysis of scattering from perfect electrically conducting surfaces. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap295.pdf.

Ding:2002:MOP

[HD02] Yun He and Chris H. Q. Ding. MPI and OpenMP paradigms on cluster of SMP architectures. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap325.pdf.

Hiraki:2002:DRU

[HIT^{+}02] Kei Hiraki, Mary Inaba, Junji Tamatsukuri, Ryutaro Kurusu, Yukichi Ikuta, Hisashi Koga, and Akira Zinzaki. Data reservoir: Utilization of multi-gigabit backbone network for data-intensive research. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

                                        /www.sc-2002.org/paperpdfs/
                                        pap.pap327.pdf.

Hacker:2002:ESP

[HNA02] Thomas J. Hacker, Brian D. Noble, and Brian D. Athey. The effects of systemic packet loss on aggregate TCP flows. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap270.pdf.

Hoschek:2002:WSD

[Hos02] Wolfgang Hoschek. The Web Service Discovery Architecture. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap161.pdf.

IEEE:2002:STI

[IEE02] IEEE, editor. SC2002: From Terabytes to Insight. Proceedings of the IEEE ACM SC 2002 Conference, November 16–22, 2002, Baltimore, MD, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1524-X. LCCN ????

Kim:2002:UDC

[KHLL02] Seung Jo Kim, Joon-Seok Hwang, Chang Sung Lee, and Sangsan Lee. Utilization of departmental computing GRID system for development of an artificial intelligent tapping inspection method, tapping sound analysis. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap167.pdf.

Kikuchi:2002:CSG

[KKN^{+}02] Hideaki Kikuchi, Rajiv K. Kalia, Aiichiro Nakano, Priya Vashishta, Hiroshi Iyetomi, Shuji Ogata, Takahisa Kouno, Fuyuki Shimojo, Kenji Tsuruta, and Subhash Saini. Collaborative simulation grid: Multiscale quantummechanical/classical atomistic simulations on distributed PC clusters in the US and Japan. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap111.pdf.

Lee:2002:MDA

[LGST02] Jason Lee, Dan Gunter, Martin Stoufer, and Brian Tierney. Monitoring data archives for grid environments. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap234.pdf.

Lee:2002:IOG

[LMN02] William Lee, Anthony Mayer, and Steven Newhouse. ICENI: An Open Grid Service Architecture implemented with Jini. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap253.pdf.

Mazzucco:2002:MMD

[MAG^{+}02] Marco Mazzucco, Asvin Ananthanarayan, Robert L. Grossman, Jorge Levera, and Gokulnath B. Rao. Merging multiple data streams on common keys over high performance networks. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap213.pdf.

Makino:2002:TSP

[MKFD02] Junichiro Makino, Eiichiro Kokubo, Toshiyuki Fukushige, and Hiroshi Daisaka. A 29.5 Tflops simulation of planetesimals in Uranus-Neptune region on GRAPE-6. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap146.pdf.

Malard:2002:DDH

[MS02] J. M. Malard and R. D. Stewart. Distributed dynamic hash tables using IBM LAPI. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap247.pdf.

Noordergraaf:2002:SSI

[NZ02] Lisa Noordergraaf and Robert Zak. SMP system interconnect instrumentation for performance analysis. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap158.pdf.

Otoo:2002:DCR

[OOS02] Ekow J. Otoo, Frank Olken, and Arie Shoshani. Disk cache replacement algorithm for storage resource managers in data grids. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap322.pdf.

Pradhan:2002:IEQ

[PcC02] Prashant Pradhan and Tzi cker Chiueh. Implementation and evaluation of a QoS-capable cluster-based IP router. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap320.pdf.

Pierce:2002:IWS

[PFY^{+}02] Marlon Pierce, Geoffrey Fox, Choonhan Youn, Steve Mock, Kurt Mueller, and Ozgur Balsoy. Interoperable Web services for computational portals. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap284.pdf.

Pike:2002:BTA

[PH02] Geoff Pike and Paul N. Hilfinger. Better tiling and array contraction for compiling scientific programs. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap260.pdf.

Parello:2002:IAA

[PTV02] David Parello, Olivier Temam, and Jean-Marie Verdun. On increasing architecture awareness in program optimizations to bridge the gap between peak and sustained processor performance – matrix-multiply revisited. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap107.pdf.

Phillips:2002:NBS

[PZKK02] James C. Phillips, Gengbin Zheng, Sameer Kumar, and Laxmikant V. Kalé. NAMD: Biomolecular simulation on thousands of processors. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap277.pdf.

Radovic:2002:ESN

[RH02] Zoran Radovic and Erik Hagersten. Efficient synchronization for nonuniform communication architectures. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap221.pdf.

Rauber:2002:LSH

[RR02] Thomas Rauber and Gudula Rünger. Library support for hierarchical multi-processor tasks. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap176.pdf.

Snavely:2002:FPM

[SCW^{+}02] Allan Snavely, Laura Carrington, Nicole Wolter, Jesus Labarta, Rosa Badia, and Avi Purkayastha. A framework for performance modeling and prediction. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap201.pdf.

Spencer:2002:EMP

[SFB^{+}02] Matthew Spencer, Renato Ferreira, Michael Beynon, Tahsin Kurc, Umit Catalyurek, Alan Sussman, and Joel Saltz. Executing multiple pipelined data analysis operations in the grid. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap258.pdf.

Sistare:2002:UHP

[SJ02] Steven J. Sistare and Christopher J. Jackson. Ultra-high performance communication with MPI and the Sun Fire(^{T M}) link interconnect. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap142.pdf.

Stamatakis:2002:APM

[SLMW02] Alexandros P. Stamatakis, Thomas Ludwig, Harald Meier, and Marty J. Wolf. Accelerating parallel maximum likelihood-based phylogenetic tree calculations using subtree equality vectors. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap113.pdf.

Sakagami:2002:TTD

[SMSY02] Hitoshi Sakagami, Hitoshi Murai, Yoshiki Seo, and Mitsuo Yokokawa. 14.9 TFLOPS threedimensional fluid simulation for fusion science with HPF on the Earth Simulator. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap147.pdf.

Shingu:2002:TGA

[STF^{+}02] Satoru Shingu, Hiroshi Takahara, Hiromitsu Fuchigami, Masayuki Yamada, Yoshinori Tsuda, Wataru Ohfuchi, Yuji Sasaki, Kazuo Kobayashi, Takashi Hagiwara, Shin ichi Habata, Mitsuo Yokokawa, Hiroyuki Itoh, and Kiyoshi Otsuka. A 26.58 Tflops global atmospheric simulation with the spectral transform method on the Earth Simulator. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap331.pdf.

Swany:2002:MRP

[SW02] Martin Swany and Rich Wolski. Multivariate resource performance forecasting in the Network Weather Service. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap292.pdf.

Schussman:2002:AVT

[SWK^{+}02] Greg Schussman, Brett Wilson, Kwok Ko, Ji Qiang, Robert Ryne, and Kwan-Liu Ma. Advanced visualization technology for terascale particle accelerator simulations. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap224.pdf.

Sterling:2002:GMP

[SZ02] Thomas L. Sterling and Hans P.

Zima. Gilgamesh: A multithreaded processor-in-memory architecture for petaflops computing. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap105.pdf.

Tapus:2002:AHT

[TCH02] Cristian Tăpuş, I-Hsin Chung, and Jeffrey K. Hollingsworth. Active Harmony: Towards automated performance tuning. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap138.pdf.

Traff:2002:IMP

[Tra02] Jesper Larsson Traff. Implementing the MPI process topology mechanism. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

                                        /www.sc-2002.org/paperpdfs/
                                        pap.pap122.pdf.

Teranishi:2002:NDM

[TRN02] Keita Teranishi, Padma Raghavan, and Esmond Ng. A new data-mapping scheme for latencytolerant distributed sparse triangular solution. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap238.pdf.

Vuduc:2002:POB

[VDY^{+}02] Richard Vuduc, James W. Demmel, Katherine A. Yelick, Shoaib Kamil, Rajesh Nishtala, and Benjamin Lee. Performance optimizations and bounds for sparse matrix-vector multiply. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap317.pdf.

Vetter:2002:APE

[VW02] Jeffrey S. Vetter and Patrick H.

Worley. Asserting performance expectations. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap275.pdf.

Vetter:2002:EPE

[VY02] Jeffrey S. Vetter and Andy Yoo.

An empirical performance evaluation of scalable scientific applications. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap222.pdf.

Wolf:2002:SPS

[WCHS02] Matthew Wolf, Zhongtang Cai, Weiyun Huang, and Karsten Schwan. SmartPointers: Personalized scientific data portals in your hand. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ???? URL http:/

/www.sc-2002.org/paperpdfs/
pap.pap304.pdf.

Bland:2002:EEI

[WDF^{+}02] P. H. Worley, T. H. Dunigan, Jr., M. R. Fahey, J. B. White III, and A. S. Bland. Early evaluation of the IBM p690. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

URL http://www.sc-2002.org/
paperpdfs/pap.pap196.pdf.

Worley:2002:SUC

[Wor02]  Patrick H. Worley.  Scaling

the unscalable: A case study on the AlphaServer SC. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap155.pdf.

Warren:2002:HDC

[WWF02] Michael S. Warren, Eric H. Weigle, and Wu-Chun Feng. Highdensity computing: A 240processor Beowulf in one cubic meter. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap210.pdf.

Ishihara:2002:TDN

[YIU^{+}02] Mitsuo Yokokawa, Ken’ichi Itakura, Atsuya Uno, Takashi Ishihara, and Yukio Kaneda. 16.4 Tflops direct numerical simulation of turbulence by Fourier spectral method on the Earth Simulator. In IEEE [IEE02], page ?? ISBN 0-7695-1524-X. LCCN ????

                                        URL http://www.sc-2002.org/
                                        paperpdfs/pap.pap273.pdf.